OpenStudy (koalagurl1):

These are the heights of Natasha and her four friends: 5 feet, 2 inches; 5 feet, 3 inches; 5 feet, 7 inches; 5 feet, 9 inches; 6 feet, 0 inches In which situation would the mean of the heights go up the most? a. You add 2 friends that are 5 feet, 8 inches. b. You add 1 friend that is 6 feet, 5 inches and 1 friend that is 5 feet, 3 inches. c. You add 1 friend that is 5 feet, 10 inches. d. You add 4 friends that are 5 feet, 9 inches.

OpenStudy (koalagurl1):

Please help! Will give medal and fan and fan testimonial!

OpenStudy (aveline):

What is the current mean?

OpenStudy (koalagurl1):

I don't get the question.

OpenStudy (photon336):

I would convert the heights to inches then find the mean of that.

OpenStudy (koalagurl1):

Oh ok thx

OpenStudy (aveline):

Start by converting the feet to inches. Right now the mean is: (62+63+67+69+72)/5

OpenStudy (koalagurl1):

hang on.

OpenStudy (koalagurl1):

The mean is 67.

OpenStudy (cutegirl1000):

The heights are, in inches, 62, 63, 67, 69, 72. The total height is 333 in. The mean height is 66.6 in. A. 333 in. + 68 in. + 68 in. = 469 in. Mean: 67 in. B. 333 in. + 77 in. + 63 in. = 473 in. Mean: 67.57 in. C. 333 in. + 70 in. = 403 in. Mean: 67.17 in. D. 333 in. + 4 * 69 in. = 609 in. Mean: 67.67 in.

OpenStudy (aveline):

Correct. Now can you apply letter A to this scenario?

OpenStudy (cutegirl1000):

So your answer is D

OpenStudy (koalagurl1):

Yeah, hang on.

OpenStudy (koalagurl1):

I need help, not the answer. @cutegirl1000

OpenStudy (cutegirl1000):

as I said before The heights are, in inches, 62, 63, 67, 69, 72. The total height is 333 in. The mean height is 66.6 in. A. 333 in. + 68 in. + 68 in. = 469 in. Mean: 67 in. B. 333 in. + 77 in. + 63 in. = 473 in. Mean: 67.57 in. C. 333 in. + 70 in. = 403 in. Mean: 67.17 in. D. 333 in. + 4 * 69 in. = 609 in. Mean: 67.67 in.

OpenStudy (koalagurl1):

Ik, but I need to do it myself.

OpenStudy (cutegirl1000):

then why are you asking others for help?

OpenStudy (aveline):

/facepalm So she can learn how to do it so she will be able to do it independently in the future. It's called learning I think..

OpenStudy (koalagurl1):

Exactly. I don't just want the answer, I want to learn how to do it so I don't have to ask again.

OpenStudy (koalagurl1):

67.28?

OpenStudy (aveline):

Sorry dc. One sec

OpenStudy (aveline):

Our original equation is (62+63+67+69+72)/5=mean=66.67 We're adding two more friends and 68 inches for each friend. Now our equation is (62+63+67+69+72+68+68)/7=mean=67.00

OpenStudy (koalagurl1):

Oh, ok.

OpenStudy (aveline):

For the next one, we're adding two friends. One is 78 inches tall, and another is 63. Now our equation is (62+63+67+69+72+78+63)/7=mean=67.71 Can you do C?

OpenStudy (aveline):

How are you doing? Do you think you've got the hang of it?